# Additive Complementary Pairs of Codes

### v2

@@ 4 Building-up construction for ACP of codes
Let $\tilde{C}_{1}$ be a linear code over ${\rm I\!F}_{q^{m}}$ generated by $\left(\begin{array}[]{ccc}0&\tilde{G}_{1}\end{array}\right)$. It is easy to see that $\tilde{C}_{1}\cap\tilde{D}_{ex_{2}}=\{0\}$ by Equation (15). Since, $\tilde{C}_{1}+\tilde{D}_{ex_{2}}\subseteq\tilde{C}_{ex_{1}}+\tilde{D}_{ex_{2}}$, $\dim_{{\rm I\!F}_{q^{m}}}(\tilde{C}_{1}+\tilde{D}_{ex_{2}})\leq\dim_{{\rm I\!F}_{q^{m}}}(\tilde{C}_{ex_{1}}+\tilde{D}_{ex_{2}})$
$\implies\dim_{{\rm I\!F}_{q^{m}}}(\tilde{C}_{1})+\dim_{{\rm I\!F}_{q^{m}}}(\tilde{D}_{ex_{2}})-\dim_{{\rm I\!F}_{q^{m}}}(\tilde{C}_{1}\cap\tilde{D}_{ex_{2}})\leq\dim_{{\rm I\!F}_{q^{m}}}(\tilde{C}_{ex_{1}})+\dim_{{\rm I\!F}_{q^{m}}}(\tilde{D}_{ex_{2}})-\dim_{{\rm I\!F}_{q^{m}}}(\tilde{C}_{ex_{1}}\cap\tilde{D}_{ex_{2}})\implies\dim_{{\rm I\!F}_{q^{m}}}(\tilde{C}_{ex_{1}}\cap\tilde{D}_{ex_{2}})\leq 1$.
Further, we choose $\lambda\in{\rm I\!F}_{q^{m}}$ such that $\det\left(\begin{array}[]{ccc}\tilde{G}_{ex_{1}}\\ \tilde{G}_{ex_{2}}\end{array}\right)=\det\left(\begin{array}[]{ccc}\lambda&P\\ 0&\tilde{G}_{1}\\ P^{\prime\top}&\tilde{G}_{2}\end{array}\right)\neq 0$. Note that there exist such $\lambda$ satisfying the condition. Then from Proposition 4.1, we have that $\left(\tilde{C}_{ex_{1}},\tilde{D}_{ex_{2}}\right)$ is an LCP of codes over ${\rm I\!F}_{q^{m}}$. ∎
We denote $C_{ex_{1}}$ be an ${\rm I\!F}_{q}$-linear additive code over ${\rm I\!F}_{q^{m}}$ corresponding to a linear code $\tilde{C}_{ex_{1}}$ over ${\rm I\!F}_{q^{m}}$ (see Equation (13)).
@@
Let $\tilde{C}$ and $\tilde{D}$ be two linear codes over ${\rm I\!F}_{q^{m}}$ such that a pair $\left(\tilde{C},\tilde{D}\right)$ is an LCP of codes (using identity $\pi$, $\sigma$ and $M$ in Proposition 4.1). Then the pair $\left(C_{ex_{1}},D_{ex_{2}}\right)$ is an ACP of codes over ${\rm I\!F}_{q^{m}}$.
@@
By Theorem 4.4 with combine Proposition 4.2, we have our results. ∎
@@
Let $\tilde{C}$ and $\tilde{D}$ be two linear codes over ${\rm I\!F}_{8}$ with generator matrices
$G_{1}=\left(\begin{array}[]{ccccccc}1&1&1&1&1&1&1\\ 1&\omega&\omega^{2}&\omega^{3}&\omega^{4}&\omega^{5}&\omega^{6}\\ 1&\omega^{6}&\omega^{5}&\omega^{4}&\omega^{3}&\omega^{2}&\omega\\ 1&\omega^{2}&\omega^{4}&\omega^{6}&\omega&\omega^{3}&\omega^{5}\\ 1&\omega^{5}&\omega^{3}&\omega&\omega^{6}&\omega^{4}&\omega^{2}\end{array}\right)$ and $G_{2}=\left(\begin{array}[]{ccccccc}1&0&1&\omega^{6}&\omega^{4}&\omega^{4}&\omega^{6}\\ 0&1&\omega^{6}&\omega^{4}&\omega^{4}&\omega^{6}&1\end{array}\right)$, respectively, where $\omega$ is a primitive root of ${\rm I\!F}_{8}$. It can be shown that $\left(\tilde{C},\tilde{D}\right)$ is an LCP of codes over ${\rm I\!F}_{8}$. By Corollary 4.5, $\left(C_{ex_{1}},D_{ex_{2}}\right)$ is an ACP of codes over ${\rm I\!F}_{8}$. However, a generator matrices of $\tilde{C}$ and $\tilde{D}$ of the form $G_{ex_{1}}=\left(\begin{array}[]{cccccccc}\omega^{5}&\omega^{6}&\omega^{3}&\omega^{4}&\omega&\omega^{2}&1&0\\ 0&1&1&1&1&1&1&1\\ 0&1&\omega&\omega^{2}&\omega^{3}&\omega^{4}&\omega^{5}&\omega^{6}\\ 0&1&\omega^{6}&\omega^{5}&\omega^{4}&\omega^{3}&\omega^{2}&\omega\\ 0&1&\omega^{2}&\omega^{4}&\omega^{6}&\omega&\omega^{3}&\omega^{5}\\ 0&1&\omega^{5}&\omega^{3}&\omega&\omega^{6}&\omega^{4}&\omega^{2}\end{array}\right)$ and $G_{ex_{2}}=\left(\begin{array}[]{cccccccc}\omega^{6}&1&0&1&\omega^{6}&\omega^{4}&\omega^{4}&\omega^{6}\\ \omega^{3}&0&1&\omega^{6}&\omega^{4}&\omega^{4}&\omega^{6}&1\end{array}\right)$.
@@
Let $C_{i}:[n,k_{i}]$ and $D_{i}:[n,n-k_{i}]$ be two linear code over ${\rm I\!F}_{q}$, for $11\leq i\leq m-$ 1. If $\alpha_{i}$’s are distinct elements of ${\rm I\!F}_{q^{m}}^{*}$ such that $\{\alpha_{0},\alpha_{1},\ldots,\alpha_{m-1}\}$ are linearly independent over ${\rm I\!F}_{q}$ and $(C_{i},D_{i})$ is an LCP for $10\leq i\leq m-$ 1, then the pair $\left(\sum\limits_{i=0}^{m-1}\alpha_{i}C_{i},\sum\limits_{i=0}^{m-1}\alpha_{i}D_{i}\right)$ is an ACP of codes over ${\rm I\!F}_{q^{m}}$.
@@
Let $C=\sum\limits_{i=0}^{m-1}\alpha_{i}C_{i}$ and $D=\sum\limits_{i=0}^{m-1}\alpha_{i}D_{i}$. Now we need to show that $\dim_{{\rm I\!F}_{q}}(C)+\dim_{{\rm I\!F}_{q}}(D)=mn$ and $C\cap D=\{0\}$. From the hypothesis, we have $\dim_{{\rm I\!F}_{q}}(C_{i})+\dim_{{\rm I\!F}_{q}}(D_{i})=n$ for $10\leq i\leq m-$ 1. Hence $\dim_{{\rm I\!F}_{q}}(C)+\dim_{{\rm I\!F}_{q}}(D)=mn$. To prove the second condition, take $x\in C\cap D$. Then $x=\sum\limits_{i=0}^{m-1}\alpha_{i}c_{i}$ and $x=\sum\limits_{i=0}^{m-1}\alpha_{i}d_{i}$ for some $c_{i}\in C_{i}$ and $d_{i}\in D_{i}$, where $10\leq i\leq m-$ 1. Therefore,
$\sum\limits_{i=0}^{m-1}\alpha_{i}c_{i}=\sum\limits_{i=0}^{m-1}\alpha_{i}d_{i}\implies\sum\limits_{i=0}^{m-1}\alpha_{i}\left(c_{i}-d_{i}\right)=0.$ (16)
Since $\{\alpha_{0},\alpha_{1},\ldots,\alpha_{m-1}\}$ are linearly independent over ${\rm I\!F}_{q}$, $c_{i}=d_{i}$, for all $i:1\leq i\leq m-1$. Further as $(C_{i},D_{i})$ is an LCP for all $i:1\leq i\leq m-1$, $c_{i}=d_{i}=0$. Hence, $C\cap D=\{0\}$. ∎
We next observe the following.
@@
Let $\tilde{C}:=[n,k]$, $\tilde{D}:=[n.n-k]$ be two linear codes over ${\rm I\!F}_{q^{m}}$ with a generator matrix $\tilde{G}$ of $\tilde{C}$ and a parity check matrix $\tilde{H}$ of $\tilde{D}$. Suppose that $\tilde{C}_{ex_{1}}$, $\tilde{D}_{ex_{2}}$ are two linear codes with a generator matrix $\tilde{G}_{ex_{1}}=\left(\begin{array}[]{cc}1&d\\ 0&\tilde{G}\end{array}\right)$ of $\tilde{C}_{ex_{1}}$ and a parity check matrix $\tilde{H}_{ex_{2}}=\left(\begin{array}[]{cc}1&c\\ 0&\tilde{H}\end{array}\right)$ of $\tilde{D}_{ex_{2}}$, where $c\in\tilde{C}_{ex_{1}}$, $d\in\tilde{C}^{\top}_{ex_{1}}$. Let $C$, $D$ be ${\rm I\!F}_{q}$-linear additive codes over ${\rm I\!F}_{q^{m}}$ corresponding to the linear codes $\tilde{C}_{ex_{1}}$ and $\tilde{D}_{ex_{2}}$, respectively over ${\rm I\!F}_{q^{m}}$ (as in Equation (13)) If the pair $(\tilde{C},\tilde{D})$ is an LCP of codes (using identity $\pi$, $\sigma$ and $M$ in Proposition 4.1), then $(C,D)$ is an ACP of codes.
@@
From Proposition 4.2, it is enough to show that $(\tilde{C}_{ex_{1}},\tilde{D}_{ex_{2}})$ is an LCP of codes. By the construction of $\tilde{C}_{ex_{1}}$ and $\tilde{D}_{ex_{2}}$, we obtain that $\tilde{C}_{ex_{1}}$ is a linear code of length $n+1$ over ${\rm I\!F}_{q^{m}}$ with dimension $k+1$ and $\tilde{D}_{ex_{2}}$ is a linear code of length $n+1$ over ${\rm I\!F}_{q^{m}}$ with dimension $n-k$. It implies that $\dim\left(\tilde{C}_{ex_{1}}\right)+\dim\left(\tilde{D}_{ex_{2}}\right)=n+1$. Now we have to show that $\tilde{C}_{ex_{1}}\cap\tilde{D}_{ex_{2}}=\{0\}$.
Let $x\in\tilde{C}_{ex_{1}}\cap\tilde{D}_{ex_{2}}$. This implies $x=\alpha\tilde{G}_{ex_{1}}=\beta G_{2}$, where $\alpha\in{\rm I\!F}_{q^{m}}^{k+1}$, $\beta\in{\rm I\!F}_{q^{m}}^{n-k}$ and $G_{2}$ is a generator matrix of $\tilde{D}_{ex_{2}}$. This gives that $\alpha\tilde{G}_{ex_{1}}\tilde{H}^{\top}_{ex_{2}}=0$ as $G_{2}\tilde{H}^{\top}_{ex_{2}}=0$. Then
$\alpha\tilde{G}_{ex_{1}}\tilde{H}^{\top}_{ex_{2}}=0\implies\alpha\left(\begin{array}[]{cc}1&d\\ 0&\tilde{G}\end{array}\right)\left(\begin{array}[]{cc}1&0\\ d^{\top}&\tilde{H}^{\top}\end{array}\right)=0\implies\alpha\left(\begin{array}[]{cc}1+dc^{\top}&0\\ 0&\tilde{G}\tilde{H}^{\top}\end{array}\right)=0.$
Since $(\tilde{C},\tilde{D})$ is an LCP of codes (using identity $\pi$, $\sigma$ and $M$ in Proposition 4.1), from Proposition 4.1, we get that $rank\left(\tilde{G}\tilde{H}^{\top}\right)=k$. Hence as $rank\left(\tilde{G}\tilde{H}^{\top}\right)=k$ and $dc^{\top}=0$, we have $\alpha=0$ i.e., $x=0$. ∎

### v3

@@ 4 Building-up construction for ACP of codes
Let $\tilde{C}_{1}$ be a linear code over ${\rm I\!F}_{q^{m}}$ generated by $\left(\begin{array}[]{ccc}0&\tilde{G}_{1}\end{array}\right)$. It is easy to see that $\tilde{\mathcal{C}}_{1}\cap\tilde{\mathcal{D}}_{ex_{2}}=\{0\}$ by Equation (13). Since, $\tilde{\mathcal{C}}_{1}+\tilde{\mathcal{D}}_{ex_{2}}\subseteq\tilde{\mathcal{C}}_{ex_{1}}+\tilde{\mathcal{D}}_{ex_{2}}$, $\dim_{{\rm I\!F}_{q^{m}}}(\tilde{\mathcal{C}}_{1}+\tilde{\mathcal{D}}_{ex_{2}})\leq\dim_{{\rm I\!F}_{q^{m}}}(\tilde{\mathcal{C}}_{ex_{1}}+\tilde{\mathcal{D}}_{ex_{2}})$
$\implies\dim_{{\rm I\!F}_{q^{m}}}(\tilde{\mathcal{C}}_{1})+\dim_{{\rm I\!F}_{q^{m}}}(\tilde{\mathcal{D}}_{ex_{2}})-\dim_{{\rm I\!F}_{q^{m}}}(\tilde{\mathcal{C}}_{1}\cap\tilde{\mathcal{D}}_{ex_{2}})\leq\dim_{{\rm I\!F}_{q^{m}}}(\tilde{\mathcal{C}}_{ex_{1}})+\dim_{{\rm I\!F}_{q^{m}}}(\tilde{\mathcal{D}}_{ex_{2}})-\dim_{{\rm I\!F}_{q^{m}}}(\tilde{\mathcal{C}}_{ex_{1}}\cap\tilde{\mathcal{D}}_{ex_{2}})\implies\dim_{{\rm I\!F}_{q^{m}}}(\tilde{\mathcal{C}}_{ex_{1}}\cap\tilde{\mathcal{D}}_{ex_{2}})\leq 1$.
Further, we choose $\lambda\in{\rm I\!F}_{q^{m}}$ such that $\det\left(\begin{array}[]{ccc}\tilde{G}_{ex_{1}}\\ \tilde{G}_{ex_{2}}\end{array}\right)=\det\left(\begin{array}[]{ccc}\lambda&P\\ 0&\tilde{G}_{1}\\ P^{\prime\top}&\tilde{G}_{2}\end{array}\right)\neq 0$. Note that there exists such $\lambda$ satisfying the condition. Then from Proposition 4.1, we have that $\left(\tilde{\mathcal{C}}_{ex_{1}},\tilde{\mathcal{D}}_{ex_{2}}\right)$ is an LCP of codes over ${\rm I\!F}_{q^{m}}$. ∎
We denote $\mathcal{C}_{ex_{1}}$ be an ${\rm I\!F}_{q}$-linear additive code over ${\rm I\!F}_{q^{m}}$ corresponding to a linear code $\tilde{\mathcal{C}}_{ex_{1}}$ over ${\rm I\!F}_{q^{m}}$ (see Equation (11)).
@@
Let $\tilde{\mathcal{C}}$ and $\tilde{\mathcal{D}}$ be two linear codes over ${\rm I\!F}_{q^{m}}$ such that a pair $\left(\tilde{\mathcal{C}},\tilde{\mathcal{D}}\right)$ is an LCP of codes (using identity $\pi$, $\sigma$ and $M$ in Proposition 4.1). Let $\mathcal{C}_{ex_{1}}$, $\mathcal{D}_{ex_{2}}$ be the codes constructed from $\tilde{\mathcal{C}}_{ex_{1}}$, $\tilde{\mathcal{D}}_{ex_{1}}$, respectively, as presented in Equation (11). Then the pair $\left(\mathcal{C}_{ex_{1}},\mathcal{D}_{ex_{2}}\right)$ is an ACP of codes over ${\rm I\!F}_{q^{m}}$.
@@
By Theorem 4.6 with combining Proposition 4.2, we have our results. ∎
@@
Let $\tilde{\mathcal{C}}$ and $\tilde{\mathcal{D}}$ be two linear codes over ${\rm I\!F}_{8^{2}}$ with generator matrices
$G_{1}=\left(\begin{array}[]{ccccccc}1&1&1&1&1&1&1\\ 1&\omega&\omega^{2}&\omega^{3}&\omega^{4}&\omega^{5}&\omega^{6}\\ 1&\omega^{6}&\omega^{5}&\omega^{4}&\omega^{3}&\omega^{2}&\omega\\ 1&\omega^{2}&\omega^{4}&\omega^{6}&\omega&\omega^{3}&\omega^{5}\\ 1&\omega^{5}&\omega^{3}&\omega&\omega^{6}&\omega^{4}&\omega^{2}\end{array}\right)$ and $G_{2}=\left(\begin{array}[]{ccccccc}1&0&1&\omega^{6}&\omega^{4}&\omega^{4}&\omega^{6}\\ 0&1&\omega^{6}&\omega^{4}&\omega^{4}&\omega^{6}&1\end{array}\right)$, respectively, where $\omega$ is $7$-th root of unity. It can be shown that $\left(\tilde{\mathcal{C}},\tilde{\mathcal{D}}\right)$ is an LCP of codes over ${\rm I\!F}_{8^{2}}$. By Theorem 4.6, $(\tilde{\mathcal{C}}_{ex_{1}},\tilde{\mathcal{D}}_{ex_{2}})$ is an LCP of codes over ${\rm I\!F}_{8^{2}}$. However, a generator matrices of $\tilde{\mathcal{C}}$ and $\tilde{\mathcal{D}}$ of the form $G_{ex_{1}}=\left(\begin{array}[]{cccccccc}\omega^{5}&\omega^{6}&\omega^{3}&\omega^{4}&\omega&\omega^{2}&1&0\\ 0&1&1&1&1&1&1&1\\ 0&1&\omega&\omega^{2}&\omega^{3}&\omega^{4}&\omega^{5}&\omega^{6}\\ 0&1&\omega^{6}&\omega^{5}&\omega^{4}&\omega^{3}&\omega^{2}&\omega\\ 0&1&\omega^{2}&\omega^{4}&\omega^{6}&\omega&\omega^{3}&\omega^{5}\\ 0&1&\omega^{5}&\omega^{3}&\omega&\omega^{6}&\omega^{4}&\omega^{2}\end{array}\right)$ and $G_{ex_{2}}=\left(\begin{array}[]{cccccccc}\omega^{6}&1&0&1&\omega^{6}&\omega^{4}&\omega^{4}&\omega^{6}\\ \omega^{3}&0&1&\omega^{6}&\omega^{4}&\omega^{4}&\omega^{6}&1\end{array}\right)$. It is easy to see that $\left(\mathcal{C}_{ex_{1}},\mathcal{D}_{ex_{2}}\right)$ is an ACP of codes over ${\rm I\!F}_{8^{2}}$ which is coincide with Corollary 4.7.
@@
Let $\mathcal{C}_{i}:[n,k_{i}]$ and $\mathcal{D}_{i}:[n,n-k_{i}]$ be two linear code over ${\rm I\!F}_{q}$, for $10\leq i\leq m-$ 1. If $\alpha_{i}$’s are distinct elements of ${\rm I\!F}_{q^{m}}^{*}$ such that $\{\alpha_{0},\alpha_{1},\ldots,\alpha_{m-1}\}$ are linearly independent over ${\rm I\!F}_{q}$ and $(\mathcal{C}_{i},\mathcal{D}_{i})$ is an LCP for $10\leq i\leq m-$ 1, then the pair $\left(\sum\limits_{i=0}^{m-1}\alpha_{i}\mathcal{C}_{i},\sum\limits_{i=0}^{m-1}\alpha_{i}\mathcal{D}_{i}\right)$ is an ACP of codes over ${\rm I\!F}_{q^{m}}$.
@@
Let $\mathcal{C}=\sum\limits_{i=0}^{m-1}\alpha_{i}\mathcal{C}_{i}$ and $D=\sum\limits_{i=0}^{m-1}\alpha_{i}\mathcal{D}_{i}$. Now we need to show that $\dim_{{\rm I\!F}_{q}}(\mathcal{C})+\dim_{{\rm I\!F}_{q}}(\mathcal{D})=mn$ and $\mathcal{C}\cap\mathcal{D}=\{0\}$. From the hypothesis, we have $\dim_{{\rm I\!F}_{q}}(\mathcal{C}_{i})+\dim_{{\rm I\!F}_{q}}(\mathcal{D}_{i})=n$ for $10\leq i\leq m-$ 1. Hence $\dim_{{\rm I\!F}_{q}}(\mathcal{C})+\dim_{{\rm I\!F}_{q}}(\mathcal{D})=mn$. To prove the second condition, take $\mathbf{x}\in\mathcal{C}\cap\mathcal{D}$. Then $\mathbf{x}=\sum\limits_{i=0}^{m-1}\alpha_{i}\mathbf{c}_{i}$ and $\mathbf{x}=\sum\limits_{i=0}^{m-1}\alpha_{i}\mathbf{d}_{i}$ for some $\mathbf{c}_{i}\in\mathcal{C}_{i}$ and $\mathbf{d}_{i}\in\mathcal{D}_{i}$, where $10\leq i\leq m-$ 1. Therefore,
$\sum\limits_{i=0}^{m-1}\alpha_{i}\mathbf{c}_{i}=\sum\limits_{i=0}^{m-1}\alpha_{i}\mathbf{d}_{i}\implies\sum\limits_{i=0}^{m-1}\alpha_{i}\left(\mathbf{c}_{i}-\mathbf{d}_{i}\right)=0.$ (14)
Since $\{\alpha_{0},\alpha_{1},\ldots,\alpha_{m-1}\}$ are linearly independent over ${\rm I\!F}_{q}$, $c_{i}=d_{i}$, for all $i:0\leq i\leq m-1$. Further as $(\mathcal{C}_{i},\mathcal{D}_{i})$ is an LCP for all $i:0\leq i\leq m-1$, $\mathbf{c}_{i}=\mathbf{d}_{i}=0$. Hence, $\mathcal{C}\cap\mathcal{D}=\{0\}$. ∎
Example 4.10.
Let $(\mathcal{C}_{1},\mathcal{D}_{1})$ be a pair of linear codes over ${\rm I\!F}_{2}$ with generator matrices $G_{1}=\left(\begin{array}[]{cccc}1&1&0&1\\ 0&1&1&1\end{array}\right)$ and $G_{2}=\left(\begin{array}[]{cccc}1&1&1&0\\ 1&0&1&1\end{array}\right)$. Also, let $(\mathcal{C}_{2},\mathcal{D}_{2})$ be an another pair of linear codes over ${\rm I\!F}_{2}$ with generator matrices $G_{3}=\left(\begin{array}[]{cccc}1&1&1&0\\ \end{array}\right)$ and $G_{4}=\left(\begin{array}[]{cccc}1&1&0&0\\ 1&0&1&0\\ 0&0&0&1\\ \end{array}\right)$. Choose, $\omega$ be an element of ${\rm I\!F}_{4}$ such that $\omega^{2}+\omega+1=0$ and $\{1,\omega\}$ forms a linearly independent set over ${\rm I\!F}_{2}$. Now, we consider $\mathcal{C}$ and $\mathcal{D}$ with generator matrices $G^{\prime}=\left(\begin{array}[]{c}G_{1}\\ \omega G_{3}\end{array}\right)$ and $G^{\prime\prime}=\left(\begin{array}[]{c}G_{2}\\ \omega G_{4}\end{array}\right)$. It is easy to see that the pair $(\mathcal{C},\mathcal{D})$ forms an ACP of codes over ${\rm I\!F}_{4}$ which coincides with the Theorem 4.9.
We next make the following proposition.
@@
Let $\tilde{\mathcal{C}}:=[n,k]$, $\tilde{\mathcal{D}}:=[n,n-k]$ be two linear codes over ${\rm I\!F}_{q^{m}}$ with a generator matrix $\tilde{G}$ of $\tilde{\mathcal{C}}$ and a parity check matrix $\tilde{H}$ of $\tilde{\mathcal{D}}$. Suppose that $\tilde{\mathcal{C}}_{ex_{1}}$, $\tilde{\mathcal{D}}_{ex_{2}}$ are two linear codes with a generator matrix $\tilde{G}_{ex_{1}}=\left(\begin{array}[]{cc}1&\mathbf{d}\\ 0&\tilde{G}\end{array}\right)$ of $\tilde{\mathcal{C}}_{ex_{1}}$ and a parity check matrix $\tilde{H}_{ex_{2}}=\left(\begin{array}[]{cc}1&\mathbf{c}\\ 0&\tilde{H}\end{array}\right)$ of $\tilde{\mathcal{D}}_{ex_{2}}$, where $\mathbf{c}\in\tilde{C}_{ex_{1}}$, $\mathbf{d}\in\tilde{\mathcal{C}}^{\top}_{ex_{1}}$. Let $\mathcal{C}$, $\mathcal{D}$ be ${\rm I\!F}_{q}$-linear additive codes over ${\rm I\!F}_{q^{m}}$ corresponding to the linear codes $\tilde{\mathcal{C}}_{ex_{1}}$ and $\tilde{\mathcal{D}}_{ex_{2}}$, respectively over ${\rm I\!F}_{q^{m}}$ (as in Equation (11)) If the pair $(\tilde{\mathcal{C}},\tilde{\mathcal{D}})$ is an LCP of codes (using identity $\pi$, $\sigma$ and $M$ in Proposition 4.1), then $(\mathcal{C},\mathcal{D})$ is an ACP of codes.
@@
From Proposition 4.2, it is enough to show that $(\tilde{\mathcal{C}}_{ex_{1}},\tilde{\mathcal{D}}_{ex_{2}})$ is an LCP of codes. By the construction of $\tilde{\mathcal{C}}_{ex_{1}}$ and $\tilde{\mathcal{D}}_{ex_{2}}$, we obtain that $\tilde{\mathcal{C}}_{ex_{1}}$ is a linear code of length $n+1$ over ${\rm I\!F}_{q^{m}}$ with dimension $k+1$ and $\tilde{\mathcal{D}}_{ex_{2}}$ is a linear code of length $n+1$ over ${\rm I\!F}_{q^{m}}$ with dimension $n-k$. It implies that $\dim\left(\tilde{\mathcal{C}}_{ex_{1}}\right)+\dim\left(\tilde{\mathcal{D}}_{ex_{2}}\right)=n+1$. Now we have to show that $\tilde{\mathcal{C}}_{ex_{1}}\cap\tilde{\mathcal{D}}_{ex_{2}}=\{0\}$.
Let $\mathbf{x}\in\tilde{\mathcal{C}}_{ex_{1}}\cap\tilde{\mathcal{D}}_{ex_{2}}$. This implies $\mathbf{x}=\alpha\tilde{G}_{ex_{1}}=\beta G_{2}$, where $\alpha\in{\rm I\!F}_{q^{m}}^{k+1}$, $\beta\in{\rm I\!F}_{q^{m}}^{n-k}$ and $G_{2}$ is a generator matrix of $\tilde{\mathcal{D}}_{ex_{2}}$. This gives that $\alpha\tilde{G}_{ex_{1}}\tilde{H}^{\top}_{ex_{2}}=0$ as $G_{2}\tilde{H}^{\top}_{ex_{2}}=0$. Then
$\alpha\tilde{G}_{ex_{1}}\tilde{H}^{\top}_{ex_{2}}=0\implies\alpha\left(\begin{array}[]{cc}1&\mathbf{d}\\ 0&\tilde{G}\end{array}\right)\left(\begin{array}[]{cc}1&0\\ \mathbf{c}^{\top}&\tilde{H}^{\top}\end{array}\right)=0\implies\alpha\left(\begin{array}[]{cc}1+\mathbf{d}\mathbf{c}^{\top}&0\\ 0&\tilde{G}\tilde{H}^{\top}\end{array}\right)=0.$
Since $(\tilde{\mathcal{C}},\tilde{\mathcal{D}})$ is an LCP of codes (using identity $\pi$, $\sigma$ and $M$ in Proposition 4.1), from Proposition 4.1, we get that $rank\left(\tilde{G}\tilde{H}^{\top}\right)=k$. Hence as $rank\left(\tilde{G}\tilde{H}^{\top}\right)=k$ and $\mathbf{d}\mathbf{c}^{\top}=0$, we have $\alpha=0$ i.e., $\mathbf{x}=0$. ∎